# AuthGlass: Enhancing Voice Authentication on Smart Glasses via Air-Bone Acoustic Features

### v2

@@ 6.4. Limited Sampling Rate
Another aspect that limit the real-world usage was the sampling rate, as our prototype samples audio at a frequency of 96 kHz, requiring relatively high energy and memory demands. In contrast, most commercial smart glasses adopt microphones with a 16 kHz sampling rate. While such a rate has little effect on Mel-spectrogram computation, it can impact sound field modeling and time-delay estimation accuracy. To verify the effect of limited sampling rates on voice authentication, we down-sampled the acoustic data to 16 kHz, and re-tested the authentication EER, with other configurations same as those in section 4.3. The results indicate that our model performs well across different sampling rates, demonstrating robustness to variations in input resolution.
@@ 7.1. Potential Applications
As our proposed method operates as a fully passive voice authentication approach on smart glasses, it has the potential to be seamlessly integrated into modern wearable devices, enabling a wide range of applications. For example, because smart glasses are convenient to use while users are engaged in other activities with their hands occupied, such as driving and cooking(Chang et al., 2018; He et al., 2015; Kim et al., 2016), AuthGlass could enable hands-free authentication for tasks like signing an urgent document or opening the smart lock, minimizing distraction and allowing users to complete the task safely. Moreover, since smart glasses with voice user interfaces are particularly beneficial for visually impaired users (Rajendran et al., 2020a; Khan and Prakash, 2017), our system can provide a secure authentication process without the risk of information leakage associated with traditional methods such as password entry. Last but not least, since voice authentication on smart glasses is closely linked to human attention, our system can be integrated with other sensors such as cameras, to enable natural interactions. For example, a user could simply look at a payment QR code and say “Pay the bill” to complete a transaction seamlessly
@@ 7.2.2. Cross-device Adaption
Though we demonstrated that vibration features and sound field features have a strong performance in liveness detection and authentication, they may face challenges in cross-device scenarios. For instance, different spatial layout of microphones will lead to different sound field signals, while the contact circumstance between vibration microphones and user skin may vary between devices, causing the vibration signal inconsistent. One of the possible solution is to leveraging cross-device adaption method by modeling the sound field across humans, which could transit the data from one types of glasses to another. What’s more, deep learning methods can be employed to enhance the generalization capability of the algorithms, as researchers have proposed unsupervised cross-domain adaptive methods(Ganin and Lempitsky, 2015)
@@ 7.2.3. Robustness against Movement Noises
Although we evaluated the performance of our system under noisy conditions, our experiments only considered scenarios in which users remained still. To support a wider range of applications, the authentication system should also be robust to noise generated by user movements. Specifically, head movements may affect vibration signals, as BC microphones can be sensitive to motion, similar to an IMU. Additionally, changes in head posture can influence the sound field features. To migrate this issue, one possible idea is to leverage additional motion-related sensors, such as IMU or gyroscope for denoising as prior works (Yan et al., 2022) has demonstrated the possibility of denoising on other motion-sensitive sensors, for example, the PPG sensor.
@@ 7.2.4. Sound Field Based Interaction on Smart Glasses
Our work demonstrates that sound-field features are related to the human speech production process, suggesting potential for broader applications. For example, Proximic (Qin et al., 2021) leveraged the time differences between a pair of earbuds to detect mouth-cover gestures. Building on this idea, sound field features could enable hands-on-mouth gesture detection during speech without relying on active acoustic sensing techniques (Guo et al., 2025a). Moreover, since the speech generation process can be influenced by the direction a user is facing, sound field features may also be used to infer the direction of the user’s speech, thereby facilitating more natural and context-aware interactions.

### v3

@@ 6.4. Limited Sampling Rate
6.4.1. Method
Another aspect that limit the real-world usage was the sampling rate, as our prototype samples audio at a frequency of 96 kHz, requiring relatively high energy and memory demands. In contrast, most commercial smart glasses adopt microphones with a 16 kHz sampling rate. While such a rate has little effect on Mel-spectrogram computation, it can impact sound field modeling and time-delay estimation accuracy. To verify the effect of limited sampling rates on voice authentication, we down-sampled the acoustic data to 16 kHz, and re-tested the authentication EER, with other configurations same as those in section 4.5.
6.4.2. Result and Discussion
As shown in Table 7, our model continues to achieve strong performance at a lower sampling rate of 16kHz. The results indicate that our model performs well across different sampling rates, demonstrating robustness to variations in input resolution.
AC AC+DMs AC+BC AC+BC+DMs
mean EER 0.0917 0.0460 0.0648 0.0302
Table 7. Averaged EER of Different Feature Combination for 16kHz Sampling Rate.
@@ 7. Discussion
7.1. Conclusion of Findings
Using our prototype, we observed that both vibration and sound-field features are closely related to human speech production. To extract these features, we applied the well-established log Mel-spectrogram method for vibration signals, and we further proposed two novel representations—time delay matrix (TDM) and energy decrease matrix (EDM)—to capture temporal variations in sound-field features. With a relatively simple authentication architecture, we demonstrated that vibration and sound-field features each improve authentication performance, and their combination yields even better results. Moreover, extensive experiments showed that the proposed AuthGlass system can effectively resist attacks while maintaining robust performance under practical conditions.
@@ 7.2. Potential Applications
As our proposed method operates as a fully passive voice authentication approach on smart glasses, it has the potential to be seamlessly integrated into modern wearable devices, enabling a wide range of applications. For example, because smart glasses are convenient to use while users are engaged in other activities with their hands occupied, such as driving and cooking(chang2018design; he2015driving; kim2016augmented), AuthGlass could enable hands-free authentication for tasks like signing an urgent document or opening the smart lock, minimizing distraction and allowing users to complete the task safely. Moreover, since smart glasses with voice user interfaces are particularly beneficial for visually impaired users (VisualImpair; khan2017design), our system can provide a secure authentication process without the risk of information leakage associated with traditional methods such as password entry. Last but not least, since voice authentication on smart glasses is closely linked to human attention, our system can be integrated with other sensors such as cameras, to enable natural interactions. For example, a user could simply look at a payment QR code and say “Pay the bill” to complete a transaction seamlessly
@@ 7.3.2. Cross-device Adaption
Though we demonstrated that vibration features and sound field features have a strong performance in liveness detection and authentication, they may face challenges in cross-device scenarios. For instance, different spatial layout of microphones will lead to different sound field signals, while the contact circumstance between vibration microphones and user skin may vary between devices, causing the vibration signal inconsistent. One of the possible solution is to leveraging cross-device adaption method by modeling the sound field across humans, which could transit the data from one types of glasses to another. What’s more, deep learning methods can be employed to enhance the generalization capability of the algorithms, as researchers have proposed unsupervised cross-domain adaptive methods(ganin2015unsupervised)
@@ 7.3.3. Robustness against Movement Noises
Although we evaluated the performance of our system under noisy conditions, our experiments only considered scenarios in which users remained still. To support a wider range of applications, the authentication system should also be robust to noise generated by user movements. Specifically, head movements may affect vibration signals, as BC microphones can be sensitive to motion, similar to an IMU. Additionally, changes in head posture can influence the sound field features. To migrate this issue, one possible idea is to leverage additional motion-related sensors, such as IMU or gyroscope for denoising as prior works (yan2022survey) has demonstrated the possibility of denoising on other motion-sensitive sensors, for example, the PPG sensor.
7.3.4. Deployment to commercial devices
As one of the key focuses of this work is the practical applicability of AuthGlass, several challenges remain for deployment on commercial devices. Although the TDM and EDM features proposed in this paper have been validated as effective, their computation still incurs a non-negligible processing time, which limits the feasibility of real-time authentication. In addition, the system relies on precise synchronization of the microphone arrays to accurately capture sound-field data, posing additional design and implementation challenges. Furthermore, there is currently no mature solution for embedding bone-conduction microphones into the nosepad of smart glasses, which may restrict both signal quality and overall system performance. Despite these limitations, advances in hardware and embedded systems present opportunities to overcome these challenges. Improved sensor integration, more efficient signal processing algorithms, and faster computation could not only resolve current hardware constraints but also enhance the real-time capability and robustness of AuthGlass, paving the way for its adoption in practical, commercial applications.
@@ 7.3.5. Sound Field Based Interaction on Smart Glasses
Our work demonstrates that sound-field features are related to the human speech production process, suggesting potential for broader applications. For example, Proximic (qin2021proximic) leveraged the time differences between a pair of earbuds to detect mouth-cover gestures. Building on this idea, sound field features could enable hands-on-mouth gesture detection during speech without relying on active acoustic sensing techniques (Echotouch). Moreover, since the speech generation process can be influenced by the direction a user is facing, sound field features may also be used to infer the direction of the user’s speech, thereby facilitating more natural and context-aware interactions.